### v1

# SpikeBottleNet: Energy Efficient Spike Neural Network Partitioning for Feature Compression in Device-Edge Co-Inference Systems

Abstract
The advent of intelligent mobile applications highlights the crucial demand for deploying powerful deep learning models on resource-constrained mobile devices. An effective solution in this context is the device-edge co-inference framework, which partitions a deep neural network between a mobile device and a nearby edge server. This approach requires balancing on-device computations and communication costs, often achieved through compressed intermediate feature transmission. Conventional deep neural network architectures require continuous data processing, leading to substantial energy consumption by edge devices. This motivates exploring binary, event-driven activations enabled by spiking neural networks (SNNs), known for their extremely energy efficiency. In this research, we propose a novel architecture named SpikeBottleNet, a significant improvement to the existing architecture by integrating SNNs. A key aspect of our investigation is the development of an intermediate feature compression technique specifically designed for SNNs. This technique leverages a split computing approach for SNNs to partition complex architectures, such as Spike ResNet50. By incorporating the power of SNNs within device-edge co-inference systems, experimental results demonstrate that our SpikeBottleNet achieves a significant bit compression ratio of up to 256x in the final convolutional layer while maintaining high classification accuracy with only a 2.5% reduction. Moreover, compared to the baseline BottleNet++ architecture, our framework reduces the transmitted feature size at earlier splitting points by 75%. Furthermore, in terms of the energy efficiency of edge devices, our methodology surpasses the baseline by a factor of up to 98, demonstrating significant enhancements in both efficiency and performance.
[…]
123
1 Introduction
In recent years, the progressive adoption of Deep Neural Networks (DNNs) in mobile technologies and Internet of Things (IoT) devices has catalyzed breakthrough performances in a wide array of intelligent applications [5]. However, DNN-based applications demand huge computational resources, rendering them impractical for direct implementation on resource-constrained mobile and edge devices. The most common deployment approach involves transmitting raw data to cloud or edge computing infrastructures for further processing and inference, followed by the resulting output being returned to the mobile device. While this approach is useful, it leads to considerable communication overhead, which causes an increase in both latency and energy usage. Furthermore, the computational power of IoT devices goes unused when data is entirely offloaded to servers. For applications involving personal data, such as health records, where privacy laws apply, local processing and anonymization are required, adding extra processing time. An alternative strategy is on-device inference, which entails the deployment of compressed DNNs on mobile devices. Nonetheless, this approach often results in compromised performance due to over-compression and is limited by the devices’ restricted resources, which in turn, increases the latency of on-device communication. These challenges have prompted the exploration of other solutions, with device-edge co-inference emerging as a viable option [19]. This method partitions a network, enabling a mobile device access to the leading part and an edge server to handle the rest of the computations. The key advantage of this approach lies in the reduction of feature tensor sizes in deeper layers of the network. By processing the earlier layers on a mobile device and then transmitting the resulting compressed features to the cloud, both latency and energy consumption associated with cloud data transmission can be substantially decreased.
Prior research on device-edge co-inference has primarily focused on traditional deep neural networks, which rely on continuous values for activations [6][28]. Although these networks deliver high performance, they are characterized by considerable computational complexity and energy demands. In contrast, distinguished as the third generation of neural network models, the brain-inspired Spiking Neural Networks (SNNs) offer a promising alternative that is extremely energy efficient when deployed in specialized neuromorphic hardware [27][2]. The adoption of binary spike signals enables SNNs to utilize low-power sparse additions as opposed to the high-power matrix multiplication operations prevalent in traditional models. SNNs also distinguish themselves by operating on binary spikes for neural communication, thereby reducing the communication overhead. Additionally, SNNs can inherently support feature compression through sparse representations and temporal coding, minimizing the need for quantization techniques aimed at shrinking feature sizes for transmission. The combination of comparable accuracy with traditional DNN [8], reduced energy consumption [29][15][26], and a high feature compression ratio makes SNNs an optimal choice for application in device-edge co-inference contexts.
This study presents a novel architecture, named SpikeBottleNet, utilizing spiking neural networks (SNNs) for image classification, emphasizing efficient feature compression within device-edge co-inference systems. The key contributions of this paper are highlighted as follows:
  • We propose a novel architecture for spike-driven feature compression in device-edge co-inference systems, which achieve comparable performance with traditional DNN based methods.
  • The experiments on CIFAR100 Datasets demonstrate that our model achieves a compression ratio of up to 256× in the final convolutional layer while maintaining an accuracy degradation of approximately 2.5%. This performance is achieved using an architecture similar to BottleNet++, but with a significant reduction in mobile energy consumption, by a factor of up to 98.
  • A key strength of SpikeBottleNet lies in its ability to facilitate the partitioning of a SNN at earlier layers. Compared to BottleNet++, SpikeBottleNet achieves a bit reduction of up to 75% at these earlier layers. This leads up to 70× reduction in energy consumption of edge devices compared to existing compression methods.
This paper continues as organized in the following manner: Section 2 summarizes an overview of related research in the field. Section 3 delves into the essential background required for comprehending the proposed work, followed by a detailed explanation of the SpikeBottleNet architecture. Section 4 outlines the experimental setup employed for our evaluations and presents a thorough account of the results obtained. The paper concludes with Section 5, which revisits the key findings and identifies potential areas for future study.
2 Background and Related Work
2.1 Deep Spiking Residual Network
In the domain of deep learning, Residual Networks (ResNet) have emerged as a prominent architecture specifically designed to tackle the degradation issue observed in increasingly deep networks. [9]. He et al. introduced the concept of identity mapping between layers through the implementation of shortcut connections. This technique significantly facilitated the training of very deep neural networks [9]. Inspired by this breakthrough in residual learning, Zheng et al. [32] established pioneering work in SNN-oriented batch normalization (BN). Their work also demonstrated the successful training of spiking ResNet-34/50 architectures using surrogate gradients within the STBP framework [30]. Extending this progression, Fang et al. [7] introduced the spike-element-wise (SEW) ResNet, employing shortcut-based integer spike merging to extend the depth of SNNs beyond 100 layers. Furthermore, Hu et al.’s Membrane Shortcut MS-ResNet introduced a novel shortcut for neuron layer membrane potentials, facilitating the training of up to 482-layer network [13]. In our study, we chose ResNet50 for its optimal balance of depth and computational efficiency, alongside its established benchmarking
These SNN based models are well-suited for implementation on neuromorphic chips, like NorthPole [23], TrueNorth [1], Loihi [24], and ROLLS [14]. These chips contain multiple interconnected neuromorphic cores (NCs) capable of running SNNs in an energy-saving manner. Unlike traditional architectures constantly processing data, these chips only activate specific neurons when receiving input spikes. This sparse operation significantly reduces power consumption, as a large portion of the chip remains idle at any given time. Compared to DNNs, the SNNs running on neuromorphic platforms can achieve up to 1000x greater power efficiency [25]. Our proposed SpikeBottleNet architecture aligns with the spike-driven paradigm, making it particularly well-suited for implementation on neuromorphic chips.
2.2 Model Splitting
Split Computing (SC) [22] is a technique that segments a DNN model into two segments: the initial segment executes on an edge device, while the subsequent portion operates on a server. This division is designed to ease the computational workload on the edge device while reducing the costs of data transmission. This is because processing delays on an edge server are typically less significant than those encountered with local processing and data transmission delays [20]. For reducing communication cost, recent research on dividing models has introduced a bottleneck feature in vision models, which results in a smaller data size than the original input sample [19][6][28][21]. Yet, these studies have not leveraged the potential of SNNs in reducing data size through model splitting and bottleneck layer, thereby overlooking an opportunity to further optimize performance.
2.3 Feature Compression
The interaction between network splitting and feature compression presents a significant challenge in the device-edge co-inference. The ideal splitting strategy seeks to minimize both on-device communication latency and feature transmission size. Feature compression has emerged as a promising approach, with numerous studies proposing various splitting and compression techniques. For instance, Li et al. in JALAD [18] employed quantization and Huffman coding for in-layer feature map compression, reducing data volume. Another study [3] presented a deep feature compression structure comprising encoder-decoder pairs, suitable for edge-cloud collaborative inference environments. BottleNet [6] introduced a bottleneck unit with a trainable lossy compressor, leveraging compressed sensing training to mitigate accuracy degradation from lossy compression. Building upon BottleNet, BottleNet++ [28] proposed an end-to-end architecture that optimizes both feature compression and encoding by integrating deep joint source-channel coding. However, BottleNet++ neglects on-device model compression, leading to excessive on-device computations at most split points. This translates to increased inference latency and energy consumption. In a more recent work [19], a novel feature compression module employing a channel attention mechanism has been introduced. This module identifies the most critical features and employs entropy encoding for eliminating statistical redundancy within the compressed data. Notably, all aforementioned research focuses on DNNs, while the potential of SNNs remain unexplored in this domain.
Figure 1: The architecture of the SpikeBottleNet
3 Proposed Method
In this section, we present the network structure for our proposed model, describing the neuron model, methods for input encoding, SNN partitioning, and feature reduction using a compression layer module, alongside our approach to training the model.
3.1 Overall Network Architecture
The architecture of our proposed model primarily comprises the design of the spike residual network and the bottleneck layer and positioning the bottleneck layer in different splitting points within the network to efficiently compress and transmit the cuboid intermediary feature tensors. As depicted in Figure 1, the feature compression module comprises an encoder and a decoder, which are designed as lightweight CNNs. These networks perform convolution operations on spikes produced by Leaky Integrate-and-Fire (LIF) neurons. This setup differs from that of BottleNet++ by substituting ReLU activation layer with LIF neurons. While Batch Normalization (BN) is commonly used in conventional DNNs for allowing stable convergence and enable deeper architectures, SNNs demand a tailored normalization technique due to their unique temporal dimension and activation dynamics. This is crucial to mitigate the problems of vanishing or exploding gradients and regulate firing rates. To address these specific requirements, we leverage threshold-dependent batch normalization (tdBN) as proposed by Zheng et al. [32].
The SNN ResNet architecture uses Residual Blocks (RBs) as its fundamental building units. Each RB is a sequence of operations: CONV-tdBN-LIF, repeated three times. An initial encoding layer first converts the input image into spike signals, which are processed through the RBs. Similar to DNN-ResNet, the number of channels doubles when the feature map size is halved. The architecture ends with average pooling and a fully connected layer to aggregate temporal information.
3.2 Spiking Neuron Model
Neural networks are composed of neurons, which transform synaptic inputs into action potential outputs. In contrast to DNNs, which ignore temporal dynamics and focuses on spatial information transmission, SNNs employ spiking neurons. These neurons simulate the behaviour of biological plausibility through the dynamics of membrane potentials and spike-based communication. The Leaky Integrate-and-Fire (LIF) model is commonly used in SNNs because it offers an optimal balance between biological plausibility and computation complexity alongside reduced energy consumption compared to DNN neurons. Our study adopts the iterative LIF model by Wu et al. [31], applied in SNNs, which uses the Euler method to convert the LIF differential equation into a recursive formula:
$V_{t}=\tau_{decay}V_{t-1}\ +\ I_{t}$ (1)
where $\tau_{decay}$ is the membrane potential decay rate, $V_{t}$ represents the membrane potential, and $I_{t}$ signifies the synaptic input at time $t$. Upon reaching a threshold $V_{th}$, a neuron fires (and $V_{t}$ resets to 0) when $V_{t}>V_{th}$. This model integrates spatial structure by considering synaptic inputs as $x_{t}=\sum_{j}{w_{j}o_{t}\left(j\right)}$ where $w_{j}$ are the synaptic weights and $o_{t}\left(j\right)$ is the binary spiking output at time $t$. The model is encapsulated by the following equations [32]:
$V_{t,n+1}=\tau_{decay}V_{t-1,n+1}\left(1-o_{t-1,n+1}\right)+x_{t,n}$ (2)
$o_{t,n+1}=\begin{cases}1,&\text{ if }V_{t,n+1}>V_{th}\\ 0,&\text{ otherwise }\end{cases}$ (3)
where $V_{t,n}$ is the neuron’s membrane potential at the $n$-th layer and time $t$, and $o_{t,n}$ represents the binary spike output.
Figure 2: Architecture of the ResNet50 and highlighting its 16 residual blocks.
3.3 Threshold-dependent batch normalization
We adopt the TDBN [32] normalization method, which operate across both spatial and temporal domains. It focuses on updating the membrane potential $u_{i}^{t}$ of a neuron at time $t$, using the TDBN to normalize a mini batch of the sequential inputs $\{I_{i}^{t}=\sum_{j=1}^{n}{w_{ij}o_{j}^{t}\ |\ t=1,\ \ldots,\ T}\}$ as follows:
$u_{i}^{t}=\tau_{mem}u_{i}^{t-1}+TDBN\ \left(I_{i}^{t}\right)$ (4)
where $TDBN$ operation on input $I_{i}^{t}$ is defined as:
$TDBN\ \left(I_{i}^{t}\right)=\lambda_{i}\frac{\alpha V_{th}(I_{i}^{t}-\mu_{c_{i}})}{\sqrt{\sigma_{c_{i}}^{2}+\epsilon}}+\beta_{i}$ (5)
Here, $\mu_{c_{i}}$ and $\sigma_{c_{i}}^{2}$ are the mean and variance for channel $i$, $\epsilon$ is a small constant for numerical stability, and $\lambda_{i}$, $\beta_{i}$, and $\alpha$ are parameters that adjust normalization, with $\alpha$ scaling inputs based on the neuron threshold $V_{th}$.
3.4 Spike Encoding
[…]
3.5 Bottleneck Unit
The bottleneck unit architecture significantly enhances data transmission efficiency between edge devices and the server, utilizing a feature reduction unit (encoder) and feature recovery unit (decoder). This design compresses data on edge devices before decompressing it in the cloud, restoring original dimensions. Both the encoder and decoder integrate three layers: a convolutional layer (or deconvolutional layer for the decoder), a threshold-dependent batch normalization (tdBN) layer, and an activation layer using LIF neurons. The encoder aims to compress data by reducing its spatial and channel dimensions, employing convolutional filters to adjust output channels and effectively manage spatial dimensions. In contrast, the decoder focuses on expanding the dimensions of the features to their original state, using a similar approach to the encoder but with an emphasis on dimension restoration. The tensor dimension changes in this process, starting from $(time\_step,batch\_size,c,w,h)$ and compressing to $(time\_step,batch\_size,c^{\prime},\ w^{\prime},\ h^{\prime})$. This involves narrowing the $c$ channels to $c^{\prime}$ and adjusting the spatial dimensions to $w^{\prime}$ and $h^{\prime}$, which is accomplished by choosing appropriate filter sizes for $c^{\prime}$ and adjusting convolution kernel strides to $\left\lfloor\frac{w}{w^{\prime}}\right\rfloor$ and $\left\lfloor\frac{h}{h^{\prime}}\right\rfloor$. This streamlined architecture ensures efficient data handling by minimizing transmission load while maintaining the integrity and usability of the data, making it particularly effective for applications requiring real-time data exchange between edge devices and cloud services.
3.6 Training Strategy
In this research, we introduce a novel architecture trainable in an end-to-end fashion. However, direct training of the entire architecture faces slow convergence. Therefore, we adopt a two-step training approach: initially training the SNN ResNet to achieve optimal accuracy [13]. The SNN ResNet consists of several layers and blocks, depicted in the Figure 2. The structure of ResNet, with its residual blocks, provides inherent breakpoints within the network that can be exploited in device edge co-inference scenarios. Next step, we deploy the bottleneck layer at these splitting points. In this step, we adopt the compression-aware training approach from [6]. The whole model is fine-tuned simultaneously by training and updating the compression module’s weights and the SNN’s parameters. Unlike the BottleNet++, which suggests a three-step approach by separately training the encoder and decoder layers in the second step while freezing the other parameters, we ignore that step to prevent slower training convergence that we observe in our experiments.
Our methodology involves placing the bottleneck unit after certain layers chosen from the underlying network, at up to $M$ potential splitting points, with $M$ being no greater than the network’s $N$ total blocks. For each point, we explore bottleneck layers that implement varying degrees of compression along channel or spatial dimensions. From those models that maintain an accuracy reduction within 3%, we select the model that requires the highest bit compression ratio. This selection process is replicated across all $M$ locations. Ultimately, we pick the best way to partition the network that minimizes edge device energy consumption and/or maximizes the bit compression ratio, in alignment with our optimization goals.
One challenge in training SNNs is the use of spike functions that can’t be differentiated. This hinders the traditional backpropagation process because the gradients can’t transmit back through the preceding layers. To overcome this, we employ a surrogate gradient approach [30], defined as:
$\frac{\partial X_{i}^{t,n}}{\partial V_{i}^{t,n}}=\frac{1}{a}sign(\left|V_{i}^{t,n}-V_{th}\right|\leq\ \frac{a}{2})$ (6)
where $a$ ensures the total gradient adds up to 1 and adjusts the curve’s steepness. This method enables differentiation through spikes, allowing our model fully end-to-end trainable.
Splitting Point RB1 RB2 RB3 RB4 RB5 RB6 RB7 RB8 RB9 RB10 RB11 RB12 RB13 RB14 RB15 RB16
BottleNet++ (B) [28] 8,192 16,384 16,384 8,192 8,192 8,192 8,192 4,096 4,096 4,096 4,096 4,096 4,096 1,024 1,024 128
SpikeBottleNet (B) 8,192 4,096 4,096 4,096 4,096 4,096 4,096 2,048 1,024 1,024 1,024 1,024 2,048 256 512 64
Bit Reduction (%) 0 75 75 50 50 50 50 50 75 75 75 75 50 75 50 50
Accuracy Drop (%) 3.45 2.33 2.13 1.40 2.94 3.70 2.34 3.63 3.90 3.13 2.64 2.85 2.61 3.76 2.53 2.52
Time Step 4 2 2 2 2 2 2 4 2 2 2 2 4 2 4 4
In the experiment, the model consistently reached the 3% accuracy degradation threshold across these splitting points. However, to investigate the trade-off between accuracy drop and compression ratio, we subsequently evaluated the model’s performance when tolerating an accuracy drop of 4%.
Table 1: Comparison of the impact of splitting points on the compressed feature size of BottleNet++ and SpikeBottleNet. Additionally, the table presents the model accuracy and bit reduction achieved by SpikeBottleNet and the Time Step used at different splitting points.
4 Evaluation
4.1 Experimental Setup
In this study, we explore a classification task that utilizes the CIFAR-100 dataset [16]. It comprises 60,000 images; each is a 32x32 pixel color image distributed across 100 categories, with 600 images in each category. This dataset is particularly suitable for deployment on edge devices since they often process low-resolution images.
For our analysis, we opted for the widely recognized ResNet50 architecture. We consider each residual block (RB) in ResNet as a potential splitting point. The ResNet50 model contains 16 residual blocks, leading to the creation of 16 distinct models. Each model is defined by placing the bottleneck unit after one of the 16 RBs. This placement results in variable degrees of reduction in both the channel and spatial dimensions across 16 different configurations. The structural design of the ResNet50 is illustrated in Figure 2.
Model Time Step Accuracy (%)
Spike ResNet50 4 76.81
2 74.69
Table 2: Accuracy of the Spike ResNet50 model at different time steps
While our spike version of ResNet50 may not surpass the current state-of-the-art performance on the CIFAR-100 dataset, the primary focus of this study lies in assessing the compression capabilities and energy efficiency achieved by our approach. As reported in the Table 2, during the training phase, the experiments achieved an accuracy of 76.81% for the Spike ResNet50 model with a time step of 4 and 74.69% with a time step of 2. To achieve a high feature compression ratio, it is essential to limit the time step to no more than 4. The cross-entropy function served as the loss function during the training process.
4.2 Compression Capability Comparison
Traditional DNNs based device-edge co-inference systems use continuous values for activations and feature compression during both training and inference phases [6]. These systems typically require high precision, often employing 32-bit floating-point (FP) representations, to preserve model accuracy. However, techniques like quantization [17] can reduce the bit-width requirements, often down to 8 bits or even lower. Conversely, our SNN based SpikeBottleNet relies on spikes, which inherently lower the bit requirements for encoding neuron states. These spikes can be efficiently encoded using just 1-bit representations, which significantly reduces the communication overhead and bitstream size. In the following, we compare the compression capability of our SpikeBottleNet with other methods, such as BottleNet++ across different splitting points.
Figure 3: Compression Capability Comparison between SpikeBottleNet and BottleNet++
The results depicted in Figure 3 reveal that, across all splitting points SpikeBottleNet consistently achieves the lowest communication overhead. This efficiency is highlighted by its achievement of up to 256x compression ratio in the final convolutional layer of the Spike ResNet50 architecture, compressing 131,072 bits (16 KB) down to 512 bits (64 Bytes), with an optimal accuracy reduction of only 2.52%. In comparison, SpikeBottleNet demonstrates a 50% reduction in the data size of 64 bytes required by the BottleNet++ (128 bytes).
A key advantage of our proposed model is its ability to facilitate earlier layer partitioning within the SNN architecture. In earlier layers, such as RB2 and RB3, SpikeBottleNet achieves a remarkable bit reduction of up to 75% compared to BottleNet++. This enhanced feature compression capability extends to layers RB9, RB10, RB11, RB12 and RB14 as well. Importantly, with the exception of RB1, SpikeBottleNet achieves at least a 50% reduction in data size across all evaluated splitting points compare to the baseline. This comparison emphasizes the SpikeBottleNet’s superior capability in feature compression for efficient communication than traditional compressors.
We study the trade-off between feature map compression and its impact on model accuracy. The Table 1 illustrates the model accuracy drop, the size of compressed features, and the bit reduction at various splitting points, alongside the time steps employed for these points. Our analysis reveals that accuracy loss remains below 3% for most of the splitting points. In some cases, however, we relaxed the accuracy threshold to 4% (RB1, RB6, RB9, RB10, and RB14) to achieve even more significant reduction in bit ratio. Notably, at RB8, no configuration could be identified that compressed features while maintaining accuracy loss of below 3%. Another key observation we noticed in our experiment is that the spatial reduction only worked for the final convolutional layer at RB16. At all other partitioning points, only channel reduction maintained the desired accuracy degradation threshold.
A key objective was to preserve a time step of 2 for compression across all layers to maximize the compression ratio. However, our findings suggest that time step 2 is only effective mostly for earlier layers, up to RB11, without a significant impact on accuracy. Beyond this point, increasing the time step to 4 becomes necessary. This suggests that the density of useful information in the feature maps increase in deeper layers. While initial layers have more redundancy that can be effectively compressed, deeper layers encode more critical information in more compact forms, thus requiring finer temporal resolution (using time step 4) to preserve this information during compression.
32 bit-FP [10]: MAC 4.6pJ, AC 0.9 pJ
Splitting Point RB1 RB2 RB3 RB4 RB5 RB6 RB7 RB8 RB9 RB10 RB11 RB12 RB13 RB14 RB15 RB16
BottleNet++ $(\mathrm{m}J)$ 0.27 0.60 0.93 1.33 1.66 1.99 2.32 2.72 3.05 3.38 3.70 4.03 4.36 4.77 5.09 5.42
SpikeBottleNet $(\mathrm{mJ})$ 0.15 0.10 0.15 0.23 0.26 0.29 0.31 0.72 0.39 0.42 0.46 0.50 0.99 0.57 1.13 1.05
$\frac{E(\text{ BottleNet }++)}{E(\text{ SpikeBottleNet })}$ 1.84 6.09 6.17 5.82 6.34 6.84 7.40 3.76 7.77 7.95 8.00 8.11 4.42 8.42 4.52 5.18
ROLLS chip[14]: 77fJ per SyOPs
SpikeBottleNet $(\mathrm{m}J)$ 0.013 0.008 0.013 0.020 0.022 0.025 0.027 0.062 0.034 0.036 0.040 0.043 0.084 0.048 0.096 0.090
$\frac{E(\text{ BottleNet }++)}{E(\text{ SpikeBottleNet })}$ 21.50 71.17 72.06 68.04 74.06 79.95 86.50 43.93 90.85 92.88 93.49 94.80 51.64 98.42 52.83 60.54
Table 3: Comparison of Mobile Energy Consumption Across Various Partition Points for Traditional 45nm Hardware and ROLLS Neuromorphic Hardware
4.3 Theoretical Energy Efficiency Estimation
Our analysis of SpikeBottleNet’s theoretical energy consumption is based on the established research [12][11]. We assume an implementation on 45nm mobile hardware, where the energy cost per MAC operation is 4.6pJ and per AC operation is 0.9pJ [10]. Our calculated energy consumption represents the edge device’s energy usage up to a specific residual block (splitting point). This analysis concentrates on the convolutional (CONV) layers within the residual blocks, which constitute a significant portion of the network’s FLOPs. The initial CONV layer converting images to spikes and the down sampling shortcuts are excluded as they deviate from the criteria for spike-based convolutions. Similarly, computations from the edge device’s encoder are omitted due to their constant operational cost.
To assess the energy consumption of SpikeBottleNet, we first calculate the cumulative synaptic operations ($SyOPs$) for each block/layer ($l$) up to a certain splitting point using the following equation:
$SyOPs(l)\ =\ fr\ \ast\ T\ \ast\ FLOPs(l)$ (7)
where, $l$ is the block or layer index in SpikeBottleNet, $fr$ represents firing rate of the input spike train reaching the block/layer, $T$ denotes simulation time step of the spiking neuron model, $FLOPs(l)$ indicate the number of floating-point operations up to layer $l$, representing the multiply-and-accumulate (MAC) operations, and $SyOPs$ signifies number of spike-based accumulate (AC) operations.
The firing rate $fr$ is defined as the probability of a neuron firing per timestep. It is estimated under a batch of random samples using the formula:
$fr\ =\ \frac{\#Spikes}{\#Neurons}\ \ast\ T$ (8)
where #$\#Spike$ and #$\#Neuron$ respectively denote the total spikes in $T$ time steps and the number of neurons in the network.
The results in Table 3 demonstrate that, on 45nm hardware, SpikeBottleNet exhibits greater power efficiency compared to BottleNet++ across all splitting points. Specifically, it achieves over 8 times at some splitting points (RB11, RB12, and RB14).
To estimate energy consumption on neuromorphic chip, we considered the ROLLS neuromorphic chip. Fabricated using a 180nm CMOS process, the ROLLS chip reportedly consumes 77fJ per SyOPs [14]. As reported in the Table 3, SpikeBottleNet exhibits up to 98 times greater power efficiency at RB14 on ROLLS hardware than BottleNet++ on conventional hardware. This finding underscores the promising future of SNN-based SpikeBottleNet. The estimated power consumption of SNNs implemented on neuromorphic hardware significantly surpasses that of DNN-based BottleNet++.
Pretraining a SNN, including the SpikeBottleNet architecutre, remains an energy efficiency challenge. The extended training time, often lasting a day even on powerful GPUs, poses a significant challenge. To address this challenge, the further development of novel algorithms and specialized hardware accelerators specifically designed for SNNs is necessary. These innovations hold the key to significantly reducing training costs and paving the way for a new era of energy-efficient neuromorphic computing.
5 Conclusions and Future Work
This work presents a novel collaborative intelligence framework, SpikeBottleNet, that tackles the latency and energy limitations hindering the deployment of deep neural networks on mobile devices. Our Spiking Neural Network (SNN) partitioning strategy significantly reduces the communication cost of transferring features from the mobile device to the cloud, while concurrently minimizing the mobile device’s energy consumption during co-inference operations at the device-edge. By exploiting the inherent binary spike signals employed by SNNs, SpikeBottleNet utilizes low-power accumulation methods instead of the high-power multiply-accumulate operations. This substitution leads to significantly minimizing the mobile device’s energy footprint. Additionally, SpikeBottleNet allows for earlier network partitioning, achieving higher compression rates than those attained by current methods. Our evaluations demonstrate that the proposed SNN-based SpikeBottleNet achieves a remarkable 256x compression ratio in the final convolutional layer. Furthermore, it translates to a concurrent 75% reduction in bit transmission at some of the earlier splitting points compared to the baseline model, with a measured accuracy degradation of only less than 3%.
Future research will focus on two key areas: communication channel modeling and input feature encoding. We aim to explore how channel characteristics like noise and bandwidth limitations impact SpikeBottleNet’s performance. This will involve developing adaptive encoding and error-correction techniques for robust spike signal transmission under varying network conditions. Additionally, we plan to investigate various input encoding methods, including attention-based mechanisms, to further enhance the network’s efficiency and adaptability.